# Type I + II Seesaw Model in light of the New Neutrino Oscillation Measurements

### v1

@@ I Introduction
The doubly and singly charged scalars that arise from the scalar sextet drive the LFV phenomenology as occurs in the canonical type II seesaw Chakrabortty:2015zpm ; Oliveira:2025kfg . Thus, we report our results in terms of their masses. In parallel to that, the doubly charged Higgs bosons that decay promptly into same-sign charged leptons have been studied in the context of the hadron collider searches Chen:2008qb ; Crivellin:2018ahj ; Chaudhuri:2013xoa ; ATLAS:2018ceg ; deMelo:2019asm ; LHeC:2020van ; Enomoto:2021fql ; Ducu:2024xxf ; Bolton:2024thn ; Zhou:2025ljo ; FCC:2025lpp ; FCC:2025uan ; FCC:2025jtd ; Chiang:2025lab . Therefore, we compare our findings with the most recent LHC results using proton–proton collision at a center-of-mass energy of 13 TeV and $36.1fb^{-1}$ of integrated luminosity ATLAS:2017xqs . Our goal is to investigate the interplay between LFV and collider physics in the framework of the 3-3-1 model.
@@ VI Acknowledgments
We thank Gabriela Hoff and Gabriela Lichtenstein for discussions. ANID - Millennium Science Initiative Program - ICN2019_044. This work is supported by Simons Foundation (Award Number:1023171-RC), FAPESP Grant 2018/25225-9, 2021/01089-1, 2023/01197-4, ICTP-SAIFR FAPESP Grants 2021/14335-0, CNPQ Grants 403521/2024-6, 408295/2021-0, 403521/2024-6, 406919/2025-9, 351851/2025-9, ANID-Millennium Science Initiative Program ICN2019_044,, and IIF-FINEP grant 213/2024. We thank IIP for the local cluster bulletcluster which was instrumental in this work.

### v2

@@ I Introduction
The doubly and singly charged scalars that arise from the scalar sextet drive the LFV phenomenology as occurs in the canonical type II seesaw Chakrabortty:2015zpm ; Oliveira:2025kfg ; Ghosh:2017pxl . Thus, we report our results in terms of their masses. In parallel to that, the doubly charged Higgs bosons that decay promptly into same-sign charged leptons have been studied in the context of the hadron collider searches Chen:2008qb ; Crivellin:2018ahj ; Chaudhuri:2013xoa ; ATLAS:2018ceg ; deMelo:2019asm ; LHeC:2020van ; Enomoto:2021fql ; Ducu:2024xxf ; Bolton:2024thn ; Zhou:2025ljo ; FCC:2025lpp ; FCC:2025uan ; FCC:2025jtd ; Chiang:2025lab . Therefore, we compare our findings with the most recent LHC results using proton–proton collision at a center-of-mass energy of 13 TeV and $36.1fb^{-1}$ of integrated luminosity ATLAS:2017xqs . Our goal is to investigate the interplay between LFV and collider physics in the framework of the 3-3-1 model.
@@ VI Acknowledgments
We thank Gabriela Hoff and Gabriela Lichtenstein for discussions. ANID - Millennium Science Initiative Program - ICN2019_044. This work is supported by Simons Foundation (Award Number:1023171-RC), FAPESP Grant 2018/25225-9, 2021/01089-1, 2023/01197-4, ICTP-SAIFR FAPESP Grants 2021/14335-0, CNPQ Grants 403521/2024-6, 408295/2021-0, 403521/2024-6, 406919/2025-9, 351851/2025-9, ANID-Millennium Science Initiative Program ICN2019_044, and IIF-FINEP grant 213/2024. The work of J.C.H. is supported by ANID FONDECYT regular No.1241685. The work of T.O. is supported by ANID FONDECYT regular No.1250343. The work of A.R. is supported by ANID FONDECYT regular No.1230110. We thank IIP for the local cluster bulletcluster which was instrumental in this work.